You are approaching a traffic signal. If the probability of encountering a red light is 0.50 and the probability of seeing a green light is 0.40, what is the probability of encountering a yellow light?

0.90

0.10

0.001

0.20

If 60% of students have an iPhone, 30% have an iMac computer, and 20% have both an iPhone and an iMac, how may students could be identified as an Apple customer having either an iPhone or iMac?

0.90

1.10

0.20

0.70

Which of the following statements is true?

If A & B are statistically independent, then P(A and B) = 0.

Statistical independence and mutually exclusive describe the same behavior.

Statistical independence is a required assumption in probability trees.

If A & B are statistically independent, then P(A|B) = P(A).

What is the probability of a score of 6 from throwing two dice?

0.167

0.028

0.138

0.111

If P(A) = 0.3 and P(B) = 0.3 where A and B are mutually exclusive events, then P(A or B) = ?

0.09

0.6

0

0.3

P(A|B) means:

Probability of event B given that event A has already occurred.

The joint probability of events A and B occurring.

Probability of event A given that event B has already occurred.

Probability of event A divided by event B.
Answer #1

1) The probabilities are independent hence P(R)+ P(G) + P(Y) = 1

given P(R) = 0.5, P(G) = 0.4 =>P(Y) = 0.1

2) If 60% of students have an iPhone, 30% have an iMac computer, and 20% have both an iPhone and an iMac, how may students could be identified as an Apple customer having either an iPhone or iMac = 0.7

P(iPhone) = 0.6 , P(iMac) = 0.3,P(iPhone n IMac) = 0.2

P(iPhone or iMac) = P(iPhone)+P(iMac)- P(iPhone n IMac) =0.6+0.3-0.2 = 0.7

3)

If A & B are statistically independent, then P(A|B) = P(A).

4) What is the probability of a score of 6 from throwing two dice?

chances are (1,5) (5,1) (2,4),(4,2) and (3,3) = 5 possible outcomes

total outcomes = 36

Probability = 5/36 = 0.138

If P(A) = 0.3 and P(B) = 0.3 where A and B are mutually exclusive events, then P(A or B) = ?

P(A or B) = P(A) + P(B) = 0.3+0.3= 0.6

P(A|B) means: Probability of event A given that event B has already occurred.